OpenStudy (anonymous):

Find an exact value. tan(7pi/12)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

-3.73205080757

OpenStudy (anonymous):

another formula for tan(x/2) = (1 - cosx)/ sinx then tan(7π/12) = (1 - cos(7π/6))/ sin(7π/6) = (1 + √3/2) / (-1/2)

OpenStudy (anonymous):

simplified -2 + √3

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

@woahitskar

OpenStudy (anonymous):

Yes, your answer would be correct.

OpenStudy (anonymous):

@lornbeach

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that isn't an answer choice though i have 2 +√3 and -2-√3 lol

OpenStudy (anonymous):

-2 + √3 would be the answer.

OpenStudy (anonymous):

It's not a choice

OpenStudy (anonymous):

It's 2 +√3. when you have -2 -√3, it becomes 2 +√3.

OpenStudy (anonymous):

Okat thats what i thought

OpenStudy (anonymous):

:)